Operations Report – April 11th

Crew 192 Operations Report 11Apr2018

SOL 3

Name of person filing report: Joe Dituri

Non-nominal systems: Structural – the main hab/EVA door.

Notes on non-nominal systems: The door swung open during the severe wind storm and knocked the top piece of the railing off. We secured the door as best we could and it swung open again. We secured it better and it’s holding for now.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 67% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas,

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 994 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133345.0

Toilet tank emptied: No. Last emptied April 9.

ATV’s Used: Red ATV, Blue ATV #2

Oil Added? No But checked and OK.

ATV Fuel Used: 5% Red, 15% Blue #2

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Spirit rover used: Yes
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 160MB at 1930.

Summary of suits and radios: Suits all charging, 2 one-piece suits used today, radios working fine.

Summary of Hab operations: Cleaned both decks, moved frig and stove to clean behind, used afternoon EVA to check up around habitat after strong winds, fixed cement step in front of rear airlock as it was not secure and almost fell apart.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: Suits inspected for safety and the Science Dome was cleaned and mopped.

Summary of RAM operations: The trailer water tank that we rolled into the RAM yesterday was examined further so we could devise the best way to clean it over the next few days.

Summary of health and safety issues: In an effort to sweep the upper deck and clean behind the refrigerator, we found a spilled box of matches on the floor near the wall directly between the fridge’s electrical system and propane stove. We removed the matches and swept the area. See Science Report for additional notes.

Questions, concerns, and requests to Mission Support: See Commander’s Report.

EVA Report – April 11th

EVA Report

EVA #4 – 11Apr2018 – Sol 3

Crew members: Joe Dituri, Andreaa Radulescu

Objectives:
Search and explore area of Route 1101 for wreckage

Operations:
0730 Begin EVA
0745 Arrive at Route 1101
0800 Progress ~halfway down route 1101 – Explored area
0900 Began to RTB
0910 Stopped at Zubrin’s Head because we though we saw some wreckage Explored for 10 min,
9940 Head back inside hab

Narrative:
Explored the Area of Route 1101. Climbed to the west most peek and was able to establish comms with the HAB from there. Also found metals cans in the shape of a P. Could be from our lost crew member…but not determinable. No other wreckage found. Suggest exploration elsewhere… Upon return we checked the state of the tank in RAM. We removed it and drained the excess water and then replaced in RAM in prep for cleaning more at some later date.

Vehicles
1. Red ATV
2. Blue #2
_______________

EVA Report

EVA #5 – 11Apr2018 – Sol 3

Crew members: Richard Blakeman, Ashok Narayanamoorthi

Objectives:
– Changed to an engineering EVA to secure hab from severe high winds
– Secured hab door, hab skirt, ladders, gear drift picked up and brought inside, broken wooden railing (and shards) brought inside
– Ensure that the hab is safe for the crew to remain in.

Operations:
1630 – 1745

Narrative:
Today’s second EVA was turned into an engineering EVA due to a strong wind storm which presented an extreme safety hazard for our crew.

Vehicles
– None

Journalist Report – April 11th

Journalist Report – 11Apr2018

Author: Victoria Varone

Today was a quieter day for Blue Crew 192. We were a little tired from all the action of the last couple of days, so a short morning EVA was followed up by lunch and then a viewing of "The Martian," a movie that the whole crew enjoys that is surprisingly similar to our situation here at the hab.

Especially now since a crazy wind storm has been raging for the last couple of hours. A couple of the crew went out on an engineering EVA to check all around the hab site and secure anything that might blow away. They found a partially broken wooden railing that had broken off and become a stake, so that and the remaining shards were recovered and brought inside.

We’re planning an Italian-style meal tonight, pasta with garlic, olive oil, and cherry tomatoes from our own GreenHab. If the wind calms down and the skies are clear enough, we might cap the night off with some stargazing.

Greenhab Report – April 11th

GreenHab Report
Name: Andreea Radulescu
Date: 11 Apr 2018

Environmental control:

Ambient with door open
Shade cloth on
Fan on (10:15~17:30)

Average temperatures: 34℃
Low 14.2℃
High 53.2℃

Hours of supplemental light: None

Changes to crops: None

Daily water usage for crops: 5 gal

Time(s) of watering for crops: Once in the morning around 10:30 am

Moringa research observations: None

Changes to research plants: No

Aquaponics: None

Narrative: Picked few tomatoes that the team will consume at dinner.

Support/supplies needed: None

Science Report – April 11th

MDRS Crew 192 Science Report (We are going to science the sh*t out of this!)

April 11, 2018 (SOL 3)

Richard Blakeman, Executive Officer

This crew has performed magnificently despite the challenges of having all the original science and engineering projects removed from the mission. The crew has pooled their individual and collective talents and shown incredible resourcefulness, creativity, imagination, and teamwork to develop multiple real-world science and engineering research and experiments.

The current list of projects that are being conducted include:

Spacesuit visor fogging study: Using off the shelf cleaning products which are used to develop mitigation strategies to solve the visor fogging issues. These are guaranteed to not scratch, damage or otherwise alter the visibility of the suits.

Hand exercises using hand relief, well-being balls: To measure comfort and hand dexterity during EVA

Crew wellness observations: Questionnaire designed to measure crew well-being during the duration of the mission

Crew weight measurements and analysis (EVA*): Pre and post weight for all personnel designed to measure body fluid loss during EVA

Crew muscle measurements: Physical measurements using tape measure to indicator of skeletal muscle glycogen reserve during mission

Ultrasonic rodent repulsion experiment: Using off the shelf, plug-in ultrasonic transmitter to observe if they deter rodent intrusions.

We noted a squeak at ~0830 today and have no confirmation but suspect there is a mouse about the house. We have also cleaned again and moved stove and frig to more thoroughly clean. What a mess behind both. Manual trap has been set IVO the squeak.

Astronomy discussions and visual observations since the main telescope is non-functional

Geology observations conducted during EVAs

EVA touch screen glove testing: We purchased several different kinds of touch pad sensitive gloves and anecdotally determining which ones are better for use with phone screen cameras.

Water contamination prevention and mitigation procedures

We have rolled the water tank from the trailer into the RAM and are going to do recurring EVAs to clean and sanitize the water tank for use with future crews to ease the requirement to make multiple trips into town for water saving fuel, vehicle wear and tear as well as cost of fuel for the society.

In honor of the achievements of Cosmonaut Yuri Gagarin, we have used a ration of our potatoes along with banana chips, apple slices and water. We heated this and combined it with yeast from the cupboard. The crew anticipates the ability to brew and enjoy a beverage made from distilled dehydrated potatoes on Yuri’s night (12 April).

The crew continues to take both still and video imagery for later analysis.

Sol Summary – April 11th

Crew 192 Sol Summary Report

Sol 3

Summary Title: Gone With the Wind

Author: Victoria Varone

Mission Status: Today was the first really hot day of our mission, a surprise in between the last few days of cool weather and wind. We continued to explore further out into the Martian terrain, exercising extra caution this time to remain hydrated and not push ourselves past our limits in hazardous conditions. The search for our missing science experiments goes on and we simultaneously improvise new science experiments to execute in their place.

Sol Activity Summary: Our day started early with our first EVA, where crew members Joe Dituri and Andreaa Radulescu went to explore another site to do a bit of hiking and geological investigation. When they returned we all had lunch together, which was followed by a brainstorming session for our next few EVAs.

Then we relaxed and watched The Martian, which was surprisingly similar to our own mission (missing crew member, same number of remaining crew members, affinity for duck tape), especially since not long after, a crazy wind storm kicked up.

Upon hearing a lot of commotion outside, we decided to have crew members Ashok Narayanamoorthi and Richard Blakeman head out on an engineering EVA prior to our scheduled evening EVA to check all around the hab. They found that a railing near the main door was broken because the front air lock door swung open and smashed railing. They made a couple of adjustments to secure the remaining piece as well as the hab door. Following that, the wind became so strong and started to present a safety hazard to the crew so the leader of the away team (Richard Blakeman) terminated the EVA.

Look Ahead Plan: We will continue to extend our range of Martian exploration as far as we possibly can, but some forecast hazardous weather has prompted us to change up some of our EVA plans. With strong winds predicted for tomorrow and Friday, we may choose to instead remain closer to the hab and work on our water tank cleanup project in the RAM and resume exploration over the weekend. Safety ALWAYS comes first.

Anomalies in Work: None

Weather: Hot and windy. The temperature reached approximately 83 degrees today and the wind storm was severe.

Crew Physical Status: All crew are physically healthy and safe.

EVA: See EVA Report.

Reports to be filed: Sol Summary Report, GreenHab Report, EVA Request, EVA Report Journalist Report, Commander’s Report, Photo Report, Science Report

Support Requested: See Commander’s Report.

Commander Report – April 11th

Mission Commander report – Crew 192 Report SOL 3

This morning I made a modification to the scheduled EVA for health and comfort of the crew. I started an EVA ~30 minutes early intentionally to avoid the heat of the day which was (late information) projected to be hotter than expected. Please do not be upset with the crew. I assumed that I had the skill and ability to make minor modifications from leading multiple missions to myriad overseas destinations with significant consequence of failure. I did so fully understanding the EVAs and other non-critical mission parameters given my role and understanding all the reasons for the established rules such as crews possibly encountering those not in SIM. I know this now to be wrong and will adjust the EVA times to account for increased temperatures by scheduling earlier in the AM and later in the PM. I had no intention of making an already tenuous situation with the society worse by my actions… I apologize for my error and will ensure it does not happen again. Please accept my humble apologies. The crew is beginning to feel a “combative / oppressive dynamic” developing between external sources on our overall team and the society into which we may be being thrust. It is not our intention to be combative or disruptive …we intend only to maximize our learning and scientific potential while here. We have an will continue to build bridges while here and work towards peace.

Please understand that EVERY piece of correspondence that leaves this sight is reviewed and approved by me and altered with my …as well as every other crew member’s input as would be the case on Mars and in any other mission. There is no one man show on this team…please consider that when questioning why my name is not on a specific report.

On a lighter note we had Martian matinee today …using the time hot midday time to catch up on a movie because the evenings have been filled with planning.

LEARNED: The heat of the day is beginning to oppress our EVAs (cut morning EVA short due to temperature) so we will schedule EVAs earlier and later to coincide with cooler points of the day. Additionally, the weather is less predictable than initially thought.

IMPRESSED: I was impressed with the skill of Blue 04 (Victoria) who has been the go between and on all email correspondence and has been dealing with an internet speed just faster than glacial as well as minimal data capacity and still manages to upload myriad required reports. BZ Blue 04!

POA&M: Tomorrow the weather is supposed to turn to exceptionally high winds. High winds have started tonight. We have done a “batten the hatches down” EVA to look around the HAB in anticipation of the pending increased wind. Many items were removed from the HAB as requested by staff on site and should be removed before winds increase. We are planning a training and cleaning day tomorrow with EVAs very close to home.

~Dituri Sends (Blue 00)